OpenStudy (anonymous):

Write an equation that has the same solution as 25 =y - 6 and has y + 7 as its right side.

OpenStudy (math&ing001):

How much do you need to add to y-6 to get y+7 ? Meaning y-6 + number = y+7

OpenStudy (math&ing001):

I just want you to try, it's ok if you get it wrong.

OpenStudy (anonymous):

Im still confused

OpenStudy (math&ing001):

- 6 +13 = 7 so y - 6 + 13 = y + 7 It's as easy as that. So now we add +13 to both sides of the original equation to get the new one : 25 = y - 6 25 + 13 = y - 6 + 13 38 = y + 7